OpenStudy (angelwings996):

exhaled air contains 16% oxygen by volume. Assuming you exhale like the average person does during strenuous exercise (that is, you exhale 1L of air per breath and you exhale 40 breaths per minute) ad assuming the dive will take 90 minutes, what volume of oxygen will you end up exhaling?

pooja195 (pooja195):

Is there some kind of formula for this? Do you know it if there is one?

OpenStudy (angelwings996):

There isn't you have to make one...like yu have to put them in fractions. I think the 90 min is 90min/1 and 16% is 16L oxy/100L air, that's all I know

syed98 (syedmohammed98):

ahh m confused...

OpenStudy (angelwings996):

I mean I can how an example one my teacher taught us. I just don't know this one

pooja195 (pooja195):

ok why not show the example we might be able to help you :)

OpenStudy (angelwings996):

okay The percent statement : The serum contains 42% antidote by mass For every 100g of serum you get 42% of antidote. So you can write it as a fraction of 42g antidote/100g serum or 100g serum/42g antidote

OpenStudy (angelwings996):

The something per something statement Serum has a density of 0.69 grams per milliliter 1 mili=0.69 grams 0.69g serum/1 mL serum or 1 mL serum/0.69g serum

pooja195 (pooja195):

wow this is really confusing maybe @Nnesha or @Marki know? i hope :(

pooja195 (pooja195):

@Abhisar help?

OpenStudy (angelwings996):

the equals or is statement 1 mili=0.034 fluid ounces 1 mL serum/0.034 fl oz serum or 0.034 fl oz/1 mL serum

OpenStudy (angelwings996):

and the last one is you need 3400 milligrams of antidote 3400 mg antidote/1 or 1/3400 mg antidote so now you put it into an equation like this:

OpenStudy (angelwings996):

since we need to find how much fl oz we need of antidote the answer needs to be fl oz ? fl oz = 0.034 fl oz/1 x 1 mL/0.69 g x 100g/42g x 1g/1000 mg x 3400 mg/1

OpenStudy (angelwings996):

you then cancel out the like terms then go the math. If it is in the numerator yu multiply and I in the denominator you divide. So it would be 3400/1000 x 100/42/0.69 x 0.034 = 0.39889579 fl oz

OpenStudy (angelwings996):

Tht's the example @pooja195 @SyedMohammed98

pooja195 (pooja195):

i am so confused on this O_o @TheRaggedyDoctor any ideas or @jagr2713

jagr2713 (jagr2713):

i dont get it to @Kainui @TheSmartOne @freckles

OpenStudy (angelwings996):

:(

TheSmartOne (thesmartone):

I honestly can't connect the question to the example either :/ @Kainui @Destinymasha @iambatman @dan815 @Mimi_x3 Can you guys understand this? Oooh! @Preetha is awesome in Chemistry. Maybe she can understand this :)

OpenStudy (dan815):

what are you confused about

OpenStudy (angelwings996):

How to set up the problem I asked. I did it in class with a partner but when I went to submit it in it was wrong.

OpenStudy (dan815):

which problem?

OpenStudy (angelwings996):

The one I asked on top that says exhaled air contains 16%

OpenStudy (dan815):

ok

OpenStudy (dan815):

exhaled air contains 16% oxygen by volume. Assuming you exhale like the average person does during strenuous exercise (that is, you exhale 1L of air per breath and you exhale 40 breaths per minute) ad assuming the dive will take 90 minutes, what volume of oxygen will you end up exhaling?

OpenStudy (angelwings996):

yes that one

OpenStudy (dan815):

http://prntscr.com/5zws4z

OpenStudy (dan815):

totalbreaths in 90 mins = 40*90=3600 1L/breath so, 3600 L of air released 16% of 3600 = 0.16*3600

TheSmartOne (thesmartone):

Thanks Dan :)

OpenStudy (angelwings996):

so 0.16*3600 will be my answer?

jagr2713 (jagr2713):

Great job Dan

OpenStudy (angelwings996):

so would the answer be 576 L oxygen ?

OpenStudy (dan815):

yes

OpenStudy (angelwings996):

Thank you !! It was right